### v1

# GuessBench: Sensemaking Multimodal Creativity in the Wild

Abstract
We propose GuessBench, a novel benchmark that evaluates Vision Language Models (VLMs) on modeling the pervasive, noisy, and pluralistic human creativity. GuessBench sources data from “Guess the Build”, an online multiplayer Minecraft minigame where one player constructs a Minecraft build given a concept (e.g., caterpillar) and others try to guess it with natural language hints, presenting a pristine testbed for sensemaking creativity in the wild with VLMs acting as guessers. We curate 1500 images from the actual gameplay and design 2000 problems spanning static and dynamic image settings, natural language hints of varying completeness, and more. Extensive experiments with six open/API VLMs and five reasoning enhancement approaches demonstrate that GuessBench presents a uniquely challenging task in creativity modeling: even the start-of-the-art GPT-4o is incorrect on 34% of instances, while we observe a huge performance gap (13.87% vs. 53.93% on average) between open and API models. When used as a resource to improve VLMs, fine-tuning on the reasoning traces for GuessBench problems improves visual perception tasks by 15.36% on average. Further analysis reveals that VLM performance in creativity sensemaking correlates with the frequency of the concept in training data, while the accuracy drops sharply for concepts in underrepresented cultural contexts and low-resource languages.
1 Introduction
Vision Language Models (VLMs) have demonstrated remarkable capabilities across perception, knowledge, and reasoning problems [61, 4, 78, 82]. From objective tasks to subjective contexts, recent works have explored creative tasks for generative models, evaluating their skills in creative generation and problem-solving across language [67, 55, 76, 51], image [13, 81, 43, 24, 60], audio [10], and video modalities [54, 23, 16, 30, 74].
We identify two key gaps in the research of creative generative modeling. 1) Creativity in the wild: while existing research holds VLMs to high artistic standards [6, 70, 77, 84], it struggles to incorporate the imperfect creativity of diverse VLM users. Their creativity is pervasive, as many of their VLM requests (e.g., polishing a photo and generating a flowchart) require creative decision making [34]; their creativity is noisy, as the average VLM user is not artistically trained and did not memorize the elite art in VLM training data [5]; their creativity is pluralistic, as different individuals could have varying interpretations for the same entity and concept [69, 17]. As such, reflecting and modeling creativity in the wild is a crucial step in aligning the creative capabilities of VLMs with diverse VLM users. 2) Sensemaking creativity: while most research focuses on generative creativity where models are tasked with generating engaging text or pretty images [49, 65, 63, 20], there is limited exploration on model capabilities in analyzing and decoding creative constructs. Quantifying and augmenting VLMs’ skills in understanding and sensemaking creativity could assist art teaching, build a reward model for creativity, and more.
Figure 1: GuessBench consists of player-constructed Minecraft builds representing real-world entities and concepts. In each problem, VLMs are required to infer what the build represents based on the provided image(s) and hint(s). We curate two settings: the static setting where only one image and hint are provided, and the dynamic setting where two sequences of progressively refined builds and corresponding hints are presented. In the dynamic setting, if the model answers correctly at any attempt, all subsequent attempts are considered correct.
To this end, we propose GuessBench, a creativity understanding dataset where VLMs guess and work out the underlying theme of a Minecraft image (Figure 1). GuessBench is sourced from “Guess the Build”, a minigame on the Hypixel Minecraft multiplayer network [^1] https://hypixel.net: Given a concept (e.g., caterpillar or oasis), one player creates a Minecraft build conveying the concept, while other players try to guess the concept with optional natural language hints. We curate 1500 images from the actual gameplay and design 2000 problems under two evaluation settings: static, where the VLM is only presented with the completed Minecraft build for one attempt, and dynamic, where the VLM is presented with a sequence of images where the build is increasingly complete for multiple attempts. We posit that GuessBench presents a pristine creativity in-the-wild setting where data is sourced from a diverse and global player community with a spectrum of language and cultural backgrounds.
We evaluate a wide range of state-of-the-art VLMs and enhancement approaches (e.g., Self-Refine [52] and Image Retrieval [87]) on GuessBench. We find that creativity sensemaking is challenging, even the state-of-the-art GPT-4o only achieves 57.8% and 66.0% in static and dynamic settings. GuessBench also introduces a quantitatively novel [41] dataset: compared to various VQA datasets [48, 22], GuessBench is more challenging, better separates model performance, and reveals novel insights about model strengths and weaknesses (Figure 2). GuessBench could also be a useful training resource: by fine-tuning VLMs on the reasoning traces in GuessBench, their performance on visual datasets improves by 15.36% on average. Further analysis reveals that VLM creativity sensemaking degrades substantially for low-resource languages, correlates with the frequency of the concept in training data, and suffers from sycophancy when facing contradictory user requests.
2 GuessBench
We propose GuessBench, a novel and challenging benchmark designed to evaluate the creative understanding capabilities of VLMs. GuessBench is based on the “Guess the Build” Minecraft minigame, where players construct a Minecraft build given a concept (e.g. caterpillar) and others try to guess it with natural language hints. We first introduce two problem settings for VLMs acting as a guesser, static and dynamic, where the Minecraft build image either stays fixed or gradually becomes more complete as the player develops their build (§2.1). To ensure the reasonableness, diversity, and difficulty of the problems, we collect images from real Minecraft gameplay scenarios and design well-crafted natural language hints with varying difficulty levels (§2.2). We then describe the evaluation metrics in §2.3 and present dataset statistics in Table 1.
2.1 Task Setting
Static Task Setting.
Given an image of a Minecraft build $b$ and a corresponding natural language hint $h$, the VLM’s goal is to decipher player creativity and identify what concept does the build represent. This task can be formally defined as: $\mathrm{VLM}(b,h)=c$, where $\mathrm{VLM}(\cdot)$ denotes the VLM, and $c$ is the textual output representing the VLM’s guessed concept (e.g. oasis or caterpillar).
Dynamic Task Setting.
Consider a sequence of Minecraft build images denoted by $\mathbf{b}_{T}=(b_{1},b_{2},\ldots,b_{T})$, where $b_{t}$ represents the build image at time step $t$. Correspondingly, we define a sequence of hints as $\mathbf{h}_{T}=(h_{1},h_{2},\ldots,h_{T})$, and a sequence of the VLM’s previous guesses as $\mathbf{c}_{T-1}=(c_{1},c_{2},\ldots,c_{T-1})$, where $\mathbf{c_{0}}$ is set as an empty sequence to indicate no previous guesses at the start. In our proposed GuessBench, both the Minecraft build images and natural language hints become progressively more complete through time, presenting a temporal and dynamic setting where VLMs need to work with incomplete information. Based on this setup, the task can be formally defined as: $\mathrm{VLM}(\mathbf{b}_{T},\mathbf{h}_{T},\mathbf{c}_{T-1})=\mathbf{c}_{T},T\geq 1$.
Statistics Value
Minecraft build sets $500$
Answer
- Unique answers/tokens $424/678$
- Maximum/Average length $4/2.1$
Static Task
Minecraft Build Images
- Total Images $500$
- Images per set $1$
- Average size (px) $1188\times 753$
- Maximum size (px) $1920\times 1080$
Hints
- Unique hints/tokens $481/120$
- Maximum/Average length $127/78.3$
Dynamic Task
Minecraft Build Images
- Total images $1500$
- Images per set $3$
- Average size (px) $1150\times 721$
- Maximum size (px) $1920\times 1080$
Hints
- Unique hints/tokens $933/125$
- Maximum/Average length $127/61.2$
Table 1: Statistics of GuessBench. The unique tokens and lengths of hints and answers are measured using the GPT-4o tokenizer.
2.2 Data Curation
Build Collection.
To collect representative visual data, we manually participate in the “Guess the Build” online multiplayer game on the Hypixel server and capture screenshots of the constructed builds. During the image selection process, we also remove low-quality builds such as those where players spell out the answer using blocks instead of building its visual representation, or where no meaningful construction is present (e.g., merely holding an apple when the target word is “apple”). In total, GuessBench comprises of 500 carefully curated build sets. For the dynamic task, we set the number of attempts $T=3$, collecting three successive Minecraft build images for each set. For the static task, we use the third-attempt image from the dynamic task (i.e., $b_{3}$) as the representative build for each set. This results in a total of 1500 images featuring diverse concepts and varying levels of completeness.
Hint Generation.
To more faithfully simulate the complete gameplay of “Guess the Build”, we draw inspiration from its progressive hint-revealing mechanism and design corresponding hints for each guessing attempt.
Specifically, suppose the ground-truth answer consists of $N$ letters (e.g., caterpillar). For the dynamic task, the first-attempt hint discloses the number of words in the answer and the number of letters in each word (e.g., _ _ _ _ _ _ _ _ _ _ _). In the second attempt, based on the previous hint, we additionally reveal $\lceil N/8\rceil$ randomly selected letters (e.g., _ _ _ _ r _ i _ _ _ _). The third-attempt hint builds on the second attempt by further providing an additional $\lceil N/4\rceil-\lceil N/8\rceil$ randomly chosen letters (e.g., c _ _ _ r _ i _ _ _ _). For the static task, we directly adopt the third-attempt hint used in the dynamic task as the sole hint for the question.
In addition to the symbolic representation, we also provide the hints in natural language to aid VLMs in better interpreting the partial information (e.g., The answer consists of 1 word. The 1st word has 11 letters. The 5th letter is ‘r’. The 7th letter is ‘i’.). Two illustrative examples for both the dynamic and static tasks are presented in Appendix H.1.
The choice of the divisors 8 and 4 serves two purposes: (1) to better mimic the hint progression in the original “Guess the Build” game, and (2) to balance the difficulty across different answer lengths, thereby enabling a more nuanced evaluation of model performance on the GuessBench.
2.3 Evaluation Metrics
To evaluate the responses generated by VLMs, we first follow the methodology proposed in MathVista [48] by employing GPT-4o to extract the predicted guess from each response (Appendix C.2), and subsequently adopt accuracy as the evaluation metric. We observe that the median and mean answer lengths across all build sets are 8 and 8.15 letters respectively. Motivated by this observation, we categorize questions with answers of length less than or equal to 8 as short answer questions, and those with longer answers as long answer questions. In addition to reporting the overall accuracy, we also report the accuracy separately for the short and long subsets, aiming to investigate whether the length of the concept word is an impact factor.
Specifically, for the three-attempt dynamic tasks, if an VLM correctly answers a build-set question in either the first or second attempt, it is exempted from answering the question again in subsequent attempts, which are automatically marked as correct.
2.4 Ethical Considerations
We take various steps to ensure the ethical compliance of GuessBench. We first inspect the Minecraft terms of service [^2] https://www.minecraft.net/en-us/terms/r2 for multiplayer servers as well as the Hypixel server’s terms of service [^3] https://hypixel.net/terms, ensuring that players consent to be viewed by a larger audience when joining multiplayer servers and the academic use and anonymized resharing of in-game content is within intended use. To protect player privacy, we apply blurring to any visible player IDs for anonymization. We manually inspect all collected images and remove any that feature hateful or offensive builds.
3 Experiment Settings
3.1 Models and Implementation
We evaluate six widely used VLMs on GuessBench, including three API and three open models. The API models include GPT-4o [61], GPT-4o-mini [61], and Gemini-2.0-Flash [71]. For the open models, we evaluate InternVL2.5-78B-MPO [78], Qwen2.5VL-72B [4], and MiniCPM-V2.6 [82].
To ensure the reproducibility, we standardize the decoding configurations across all VLMs by setting the temperature to 0.0 and top_p to 1.0, or by disabling sampling via setting do_sample to False. The specific prompts and hyperparameters used for each VLM are detailed in Appendix C. All experiments are conducted using eight NVIDIA A100-SXM4-40GB GPUs.
3.2 Reasoning Approaches
The models we tested employ chain-of-thought (CoT) reasoning by default, while we further explore several alternative reasoning strategies specifically for the strongest model GPT-4o. The results are reported in Table 2, and the corresponding instructions for each approach are provided in the Appendix C.3.
In the w/o CoT configuration, the prompt explicitly instructs the model to generate a direct answer without any intermediate reasoning steps. We explore this setup to investigate whether the intermediate reasoning steps are helpful in the task of creativity sensemaking.
In the One-shot setting, we provide a single demonstration. For static tasks, this demonstration includes an image of a Minecraft build, a hint, and the corresponding answer. For dynamic tasks, the demonstration consists of the current attempt along with all previous attempts, each accompanied by its associated image, hint, and answer.
The Self-Consistency [79] generates three independent responses and adopts the final prediction through majority voting. If all three responses differ, the last guess is selected as the final answer.
The Self-Refine [52] approach prompts the model to evaluate its own response. If it deems the initial prediction correct, it outputs the answer directly. Otherwise, it reconsiders its response up to two additional times. If the model is still unable to affirm the correctness of any prediction, the last guess is adopted as the final answer.
Building on this, and tailored to the unique characteristics of our task, we further propose an extension of the Self-Refine method that incorporates Image Retrieval [87] to better support the model’s self-evaluation. During self-evaluation, the model uses its current guess as a query to retrieve the top image result from Google Images [^4] https://www.google.com/imghp.. This retrieved image is then incorporated into the model’s reasoning to reassess the validity of its prediction.
4 Results
Model Static Dynamic #1 Dynamic #2 Dynamic #3
Short Long All Short Long All Short Long All Short Long All
API Models
GPT-4o 58.7 56.8 57.8 10.8 12.0 11.4 40.2 44.4 42.2 65.6 66.4 66.0
Gemini-2.0-Flash 47.5 44.0 45.8 8.5 7.5 8.0 29.3 25.7 27.6 52.9 50.6 51.8
GPT-4o-mini 40.9 32.4 36.8 7.3 7.1 7.2 29.0 23.7 26.4 49.8 37.8 44.0
Open Models
InternVL2.5-MPO 25.5 18.7 22.2 3.1 2.1 2.6 12.0 7.9 10.0 23.9 14.1 19.2
Qwen2.5VL-72B 15.4 8.3 12.0 2.3 0.8 1.6 8.5 6.2 7.4 18.1 9.5 14.0
MiniCPM-V2.6 12.4 6.2 9.4 1.9 1.2 1.6 4.6 3.3 4.0 10.8 5.8 8.4
Single-Modality Inputs (GPT-4o)
Visual Input Only 23.2 22.0 22.6 6.9 6.2 6.6 17.0 14.5 15.8 30.1 26.6 28.4
Hint Only 17.0 18.3 17.6 0.4 1.7 1.0 7.7 12.0 9.8 20.8 28.2 24.4
Augmented GPT-4o
w/o CoT 59.1 58.1 58.6 11.6 12.0 11.8 42.1 44.4 43.2 68.0 67.6 67.8
One-shot 59.8 53.1 56.6 10.4 11.2 10.8 42.5 42.7 42.6 70.7 64.7 67.8
Self-Consistency 59.1 55.2 57.2 11.2 13.3 12.2 39.4 42.3 40.8 70.7 65.6 68.2
Self-Refine 61.0 58.1 59.6 9.3 10.4 9.8 43.2 46.1 44.6 70.3 67.6 69.0
Image Retrieval 58.3 53.9 56.2 9.3 12.0 10.6 42.1 42.3 42.2 67.6 65.6 66.6
Table 2: Evaluation results on GuessBench. Bold values indicate the best performance within each category. All denotes the overall accuracy for each task. One model name is abbreviated due to space limits. InternVL2.5-MPO: InternVL2.5-78B-MPO. The results demonstrate the challenging nature of GuessBench and reveal a clear performance disparity between open and API models.
Table 2 presents the performance of six open and API models on GuessBench, from which we draw several key observations.
First, GuessBench is highly challenging. The best-performing model GPT-4o achieves only 57.8% accuracy on static tasks and 66.0% accuracy on the final attempt of dynamic tasks. This indicates that in at least 34.0% of cases, even the strongest VLM wasn’t able to decipher user creativity.
Second, we observe a significant performance gap between open and API models. GPT-4o outperforms the best open-source model by 160% in static tasks and by 244% in the third attempt of dynamic tasks. This disparity indicates that, although open models demonstrate competitive or even superior performance to API models on earlier benchmarks [4, 78], their performance degrades substantially on GuessBench’s novel tasks. This suggests that previous benchmark success may stem from memorization of similar data and tasks, rather than a genuine and overall improvement of visual perception and reasoning. Consequently, open models exhibit limited creative and conceptual understanding.
Third, API models benefit considerably from the iterative refinement in the dynamic setting. In contrast, open models show minimal improvement, and in some cases, performance even declines. For the third attempt in the dynamic task, models are additionally provided with less complete Minecraft builds and imperfect hints from the previous two attempts. Under this setting, the average accuracy of API models increases by 15.24% compared to their static task performance. However, open models exhibit little to no improvement, and the performance of InternVL2.5-78B-MPO even decreases by 13.51%. This indicates that open models struggle to understand the incomplete builds in GuessBench, further revealing their limited ability to generalize to unfamiliar tasks.
Fourth, none of the reasoning strategies tested leads to substantial performance improvements. Among them, Self-Refine proves to be most effective, while Image Retrieval performs poorly. Self-Refine increases accuracy by only 3.11% on static tasks and 4.55% on the third attempt of dynamic tasks. This outcome suggests that the challenges posed by GuessBench cannot be adequately addressed through existing strategies to enhance VLMs. Interestingly, Image Retrieval causes a 2.77% decrease in static task accuracy and only yields a 0.91% increase in dynamic task attempt 3. Our in-depth case study reveals that some questions in GuessBench have multiple valid answers, each corresponding to a different visual representation. When a model’s initial guess is correct, the retrieved image may reflect an alternative but still valid construction. However, due to differences between the original and retrieved images, the model may become confused and abandon its correct initial guess (Appendix H.2). This highlights a current limitation in the model’s ability to handle many-to-one mappings between images and textual answers, particularly in scenarios requiring creative visual perception. It also indicates that while most existing VLM enhancement methods on evaluated on common domains such as VQA and math, progress might not be generalizable to the wide spectrum of other VLM uses such as comprehending user creativity.
Lastly, GuessBench is fundamentally a cross-modal benchmark, where relying solely on text or image input is insufficient. On static tasks, GPT-4o achieves only 22.6% accuracy when given image-only input and 17.6% accuracy with text-only input. These are both substantially lower than the 57.8% accuracy achieved with multimodal input. A similar pattern is observed in dynamic tasks. This demonstrates that a single modality alone is inadequate for solving the tasks in GuessBench, highlighting the necessity of effective multimodal integration for successful task completion.
5 Analysis
5.1 Quantitative Evaluation of GuessBench via AutoBencher Metrics
Figure 2: The AutoBencher scores of GuessBench, MathVista, and HallusionBench, higher is better. GuessBench demonstrates higher levels of difficulty, separability, and novelty compared to the other two benchmarks.
We adopt the evaluation metrics proposed by the AutoBencher [41] framework to quantitatively assess the quality of our dataset. Specifically, we compute three key metrics: Difficulty, Separability, and Novelty. The Difficulty of a benchmark is defined as the lowest error rate achieved by any model, reflecting the overall challenge posed by the task. Separability measures the degree to which models can be distinguished based on their performance, calculated as the mean absolute deviation of accuracies across models. Novelty quantifies how distinct the benchmark is from existing ones, computed as one minus the rank correlation between the accuracy vector on the current dataset and the most similar predicted accuracy vector derived from existing datasets. These metrics provide a quantitative measurement of dataset and task quality.
To compute these metrics, we evaluate five vision-language models: two API models (GPT-4o [61] and Gemini-2.0-Flash [71]) and three open models (InternVL2.5-78B-MPO [78], Qwen2.5VL-72B [4], and MiniCPM-V2.6 [82]). We benchmark GuessBench alongside MathVista [48] and HallusionBench [22], with the results summarized in Figure 2. The detailed accuracy of all models on all involved benchmarks is presented in Appendix C.6.
Difficulty.
GuessBench achieves a difficulty score of 0.422, which is 80.34% higher than MathVista and 58.05% higher than HallusionBench. This significant increase indicates the heightened complexity and challenge posed by GuessBench.
Separability.
GuessBench obtains a separability score of 0.183, which is 2.86 times that of MathVista and 7.96 times that of HallusionBench. This suggests that GuessBench is substantially more effective at differentiating the performance of various models in understanding creativity in the wild, revealing the strengths and weaknesses of models that were unclear with previous datasets.
Novelty.
We assess novelty by considering MMBench V1.1 Test (EN) [45] and AI2D [33] as prior datasets. Based on this metric, GuessBench achieves higher novelty score than both MathVista and HallusionBench. Moreover, the model-specific accuracy scores on GuessBench reveal a clear performance gap between open-source and proprietary models in addressing creative reasoning tasks. This further underscores the novelty and significance of our dataset and benchmark task in evaluating real-world multimodal understanding.
5.2 Evaluating Multilingual Robustness in Creative Tasks
Figure 3: Performance of GPT-4o across eight languages with varying levels of resource availability. In the chart, r denotes the Pearson correlation coefficient. For low-resource languages, GPT-4o exhibits a noticeable decline in performance across four different settings. The performance shows a strong correlation with how widely each language is spoken.
To investigate whether VLMs exhibit performance disparities across different linguistic environments, we conduct a multilingual evaluation based on the tasks categorized as Visual Input Only in Table 2. For each selected task, we provide identical prompts translated into various languages and translate the corresponding answers into the same target languages. GPT-4o is then tasked with reasoning under different language settings of the same problem.
To quantify language prevalence, we refer to the number of speakers (in millions) for each language as reported by Wikipedia, and visualize the relationship in Figure 3. We observe that for high-resource languages such as English and Chinese, GPT-4o achieves significantly higher accuracy across all three attempts on both static and dynamic tasks, compared to languages with fewer speakers. Across all tasks, the Pearson correlation coefficient between the number of speakers and GPT-4o’s accuracy is at least 0.810, indicating a strong positive correlation.
These findings suggest that the creative sensemaking capabilities of VLMs are substantially reduced in low-resource language settings, revealing a critical gap in current multilingual generalization and raising concerns about the equity of model performance when serving diverse language speakers.
5.3 Impact of Concept Frequency
To investigate the impact of concept frequency on model performance, we adopt the Infini-gram [44] approach and use Dolma-v1.7 (2.6T tokens) as our corpus to obtain concept frequency statistics for each problem. We then evaluate the accuracy of GPT-4o under the static task setting. The accuracy for each bin represents the average accuracy of all samples within that range. As shown in Figure 4, we visualize the results using two methods: on the left, the x-axis represents the top-k percentage of answers sorted in ascending order of frequency; on the right, the x-axis directly reflects the raw word frequency in ascending order. We compute the Pearson correlation coefficient between accuracy and average word frequency across bins, yielding values of 0.852 and 0.770 respectively, indicating a strong positive correlation. These results suggest that when GPT-4o attempts to understand creativity in the wild, it is more likely to succeed with higher-frequency concepts, revealing limitations in its sensemaking capabilities for long-tail entities and concepts underrepresented in the training data.
Figure 4: GPT-4o performance across different levels of concept frequency. Bins containing fewer than three answers are omitted and shown as blank. Darker bin colors indicate higher accuracy values. In the chart, r denotes the Pearson correlation coefficient. Both plots reveal a consistent trend: GPT-4o’s accuracy increases with concept frequency.
5.4 Effect of Multi-View Inputs
Figure 5: GPT-4o performance under Single-View and Multi-View settings. In three successive attempts of the Dynamic task, having Minecraft build images from different angles affects GPT-4o’s performance on both short-answer and long-answer questions. However, as the number of attempts increases, the impact gradually diminishes, while the overall performance remains largely consistent.
To explore whether providing VLMs with additional physical views of the same Minecraft build enhances their creative sensemaking, we evaluate GPT-4o under the Dynamic task setting. Specifically, we sample 101 sets of Minecraft build images, and for each attempt within the Dynamic task, we supplement the original input with two additional images captured from different viewpoints of the same build.
As shown in Figure 5, an intriguing pattern emerges: when analyzing short-answer and long-answer questions individually, additional viewpoints have some influence on GPT-4o’s accuracy, but this influence decreases markedly with more attempts. In contrast, when considering overall accuracy across both question types, the impact of additional viewpoints becomes minimal or even negligible.
These findings suggest that merely offering more visual angles of a build does not substantially aid in identifying what the build represents, echoing VLMs’ limitations in multi-view reasoning [85]. This further highlights the challenging nature of GuessBench, as it requires VLMs to engage in deeper creative understanding, association, and reasoning rather than relying solely on expanded visual input.
5.5 Impact of GuessBench Fine-Tuning on Related Tasks
To investigate whether creativity sensemaking benefits VLMs on other tasks, we conduct a series of transfer learning experiments using the Qwen2-VL-7B model. We evaluate performance under three distinct settings: without tuning, synthetic tuning, and mixed tuning, across three benchmarks: MathVista [48], MultiChartQA [88], and BLINK [18]. Detailed results are shown in Figure 6.
For the synthetic tuning setting, we collect all instances correctly answered by GPT-4o on GuessBench under the static task setting, yielding 289 data points. We then fine-tune Qwen2-VL-7B using these questions and GPT-4o response pairs. For the mixed tuning setting, we augment the synthetic tuning data with an additional 289 examples sampled from tasks correctly answered by GPT-4o in the target benchmarks, resulting in a total of 578 training examples. For evaluation, we exclude these training samples from each benchmark and randomly select 289 held-out examples per benchmark as the test set. All fine-tuning hyperparameters are kept consistent across conditions, as detailed in Appendix C.5.
Our findings show that synthetic tuning leads to noticeable performance improvements on MultiChartQA and BLINK, although performance on MathVista declines. Under mixed tuning, accuracy improves further on MultiChartQA and BLINK, with particularly substantial gains on MultiChartQA, while MathVista performance continues to decline.
Figure 6: The performance of various fine-tuning methods. Fine-tuning improves performance on both MultiChartQA and BLINK, whereas it reduces performance on MathVista.
Since GuessBench is designed to emphasize perceptual and interpretive capabilities, fine-tuning on it significantly boosts performance on perception-driven tasks like BLINK and, to a lesser extent, MultiChartQA. In contrast, it appears less beneficial and potentially detrimental for tasks that demand precise logical or mathematical reasoning, such as those in MathVista. Although both synthetic and mixed tuning promote CoT reasoning in Qwen2-VL-7B, the resulting gains are largely confined to tasks requiring high-level visual understanding and flexible sensemaking. These results suggest that GuessBench effectively enhances a model’s interpretive competence, making it better suited for tasks where nuanced perception and contextual inference are critical.
6 Related Work
VLM Evaluation.
VLMs are increasingly popular and demonstrate great performance across a wide range of vision-language tasks. As such, evaluating their capabilities has emerged as a critical research problem. Existing benchmarks predominantly focus on objective tasks, including: (1) General Question Answering, such as MMBench [45], SEED-Bench [38], and MMMU [83]; (2) Optical Character Recognition (OCR), such as OCR-VQA [56] and OCRBench [46]; (3) Graphic Understanding, including AI2D [33], CharXiv [80], and MultiChartQA [88]; and (4) Mathematics, such as MathVista [48], MathVision [75], and MathV360K [68]. However, current VLM benchmarks fall short in systematically exploring the subjective capabilities of these models. In this work, we propose to evaluate VLMs through tasks that assess understanding and sensemaking creativity through GuessBench, offering a lens into the models’ performance on more subjective dimensions.
Creativity in Generative Models.
Prior work on the creativity of generative models has primarily focused on elite-level creativity, such as high-quality writing [6, 72, 49, 37, 35, 21, 8, 3, 53, 9], scientific discovery [32, 25], image synthesis [26, 66, 31, 27, 15, 47], problem solving [73, 2, 1, 19, 58], code generation [50, 36], combinational creativity [86, 63, 57, 14, 40], and embodied reasoning [62, 81, 12, 28, 7, 39, 64, 42]. However, prior research has largely overlooked creativity at the level of the general public, which can be described as creativity in the wild. While such creativity may be imperfect, it is widespread and highly personalized. To address this gap, we propose GuessBench as a proxy for quantifying the ability of VLMs to understand the creativity in the wild.
7 Conclusion
We present GuessBench, a novel benchmark designed to evaluate VLMs’ capacity for creativity sensemaking in the wild. By leveraging gameplay data from the Minecraft minigame “Guess the Build”, our dataset captures the pervasive, noisy, and pluralistic nature of real-world human creativity, offering a unique and challenging testbed for VLMs acting as guessers. Through comprehensive experiments, we demonstrate that state-of-the-art models such as GPT-4o struggle with dynamic reasoning and decoding linguistically/culturally underrepresented contexts. Our findings underscore the limitations of existing VLMs in modeling diverse, imperfect forms of creativity and highlight the importance of training on realistic, user-generated data. GuessBench not only reveals nuanced gaps in model performance but also shows potential for enhancing VLM capabilities through fine-tuning on reasoning traces for creativity sensemaking. We envision GuessBench as a valuable resource in aligning VLMs more closely with the complex and inclusive landscape of human creative expression.
Appendix A Limitations
There are two primary limitations. First, due to disparities in technological development across countries and regions, the majority of players in the “Guess the Build” game on the Hypixel server are from more developed areas. As a result, the data we collect lacks balanced geographical representation at the global level. Second, in our dynamic task setting, current VLMs do not yet support directly answering questions based on video input. Therefore, we simulate the dynamic process using three static images accompanied by corresponding hints. We anticipate that both limitations can be addressed in the future as global technological infrastructure improves and any-to-any models become available.
Appendix B Ethics Statement
Due to the presence of social and cultural biases [59], VLMs tend to exhibit higher accuracy when predicting certain concepts over others, and their interpretations often align more closely with Western cultural contexts. For example, as shown in Figure 7, the Minecraft build represents the logo of China Post. While players familiar with this symbol can easily identify it, GPT-4o describes it as “fast food.” This example illustrates that VLMs exhibit a clear cultural bias in their understanding of creative content.
Figure 7: A Minecraft build depicting the China Post logo.
Appendix C Experiment Details
C.1 Instructions
Table 3 presents the default instructions used in GuessBench. For the static task, where only a single guess is allowed, we use the Initial Guess Instruction exclusively. For the dynamic task, the Initial Guess Instruction is used for the first attempt. If the response is incorrect, subsequent attempts incorporate the model’s previous guess to construct the Guess Refinement Instruction.
Category Detail
System You are an AI designed to play the “Guess the Build” game from the Hypixel server in Minecraft. Your task is to analyze a player’s build and accurately guess what it represents. The build is constructed using Minecraft blocks, and it can depict objects, animals, structures, abstract concepts, or any other recognizable entity. Consider common themes, contextual clues, and typical constructions in Minecraft. If a build is unclear, make educated guesses based on possible interpretations. Your goal is to guess as accurately and efficiently as possible. Avoid overly broad answers—be specific when possible. If multiple plausible answers exist, prioritize the most likely one based on common builds in the game.
Initial Guess Look at the provided image of a Minecraft build and the corresponding hint. Based on its structure, shape, materials, and the given hint, determine what it represents. Consider common objects, animals, buildings, abstract concepts, or any other recognizable entities typically built in Minecraft. You can use chain of thought reasoning to arrive at the best possible guess.
Guess Refinement Your previous guess is: previous_guess. This guess is incorrect. Now, you will be provided with a more complete version of the build along with additional hints. Analyze the new information carefully and make a revised guess based on the updated build and hints. You can use chain of thought reasoning to improve your guess.
Table 3: Illustrations of Instructions. The table presents the default instructions used for evaluating GuessBench. During evaluation, either the Initial Guess Instruction or the Guess Refinement Instruction is concatenated with the corresponding hint and then provided as input to the VLM.
C.2 Answer Extraction Prompt
Table 4 presents the instruction provided to GPT-4o for extracting the predicted guess generated by the VLMs.
Instruction
I will give you a language model’s response in a ‘Guess the Build’ game. Identify and extract the model’s guessed answer. Output it in the format: ‘Answer: [extracted answer]’. If no answer is given in the response, output ‘No Answering’.
Language model’s response:
Table 4: Illustrations of the instructions used for guess extraction. The response from the VLM is appended directly after these instructions, and the combined input is then fed into GPT-4o to extract the predicted guess.
C.3 Prompt for Different Reasoning Approaches
Table 5 presents the core instructions for the different reasoning approaches. For the Self-Consistency method, no specific instruction is provided; instead, the VLM is prompted three times, and the final answer is determined via majority voting.
Approach Detail
Without CoT You are not allowed to use chain-of-thought reasoning. You must output the final answer directly.
One-shot I will give you an example to help you guess the build. Here is an example:
<Image>
Hint_3: The answer format is as follows:
aalaaaasaaaaaaaaaaaak
_ _ _ _ _aa_ _aa_ _ _ _
This means the answer consists of 3 words. The 1st word has 5 letters. The 2nd word has 2 letters. The 3rd word has 4 letters. The 1st word’s 2nd letter is l. The 1st word’s 5th letter is s. The 3rd word’s 4th letter is k.
Answer: The build represents Glass of Milk.
Self-Refine Here is your response: previousresponse Please review your answer to check if it aligns with the given hint. If it does, you MUST ONLY output “\$Well done!\$" Otherwise, provide your improved guess.
Image Retrieval Previously, your guess was: previous_guess Now, I have searched for a real image of previous_guess from Google Images and attached it above. Please compare the provided Minecraft build with the real image of your guess. If they match, you MUST ONLY output: “\$Well done!\$". Otherwise, provide your improved guess.
Table 5: Illustrations of the instructions used for different reasoning approaches. For each reasoning approach, we present the core instruction to provide a more intuitive and accessible understanding.
C.4 Model Hyperparameters
Table 6 presents the detailed parameter configurations for the models discussed in (§3.1). Parameters not explicitly specified follow their default settings. Notably, for the Self-Consistency experiments described in (§3.2), we allow sampling and do not manually set the temperature or top_p parameters, instead relying on their default values. This design choice is intended to promote response diversity during generation, thereby ensuring that the results obtained through majority voting are meaningful.
Model Generation Setup
GPT-4o model = gpt-4o-2024-11-20, temperature = 0.0, top_p=1.0, max_tokens = 1000
Gemini-2.0-Flash model = gemini-2.0-flash-001, temperature = 0.0, top_p=1.0, max_output_tokens=1000
GPT-4o-mini model = gpt-4o-mini-2024-07-18, temperature = 0.0, top_p=1.0, max_tokens = 1000
InternVL2.5-MPO max_new_tokens = 1000, do_sample = False
Qwen2.5VL-72B max_new_tokens = 1000, do_sample = False
MiniCPM-V2.6 max_new_tokens = 1000, sampling = False
Table 6: Generating parameters for VLMs.
C.5 Implementation Details for Fine-tuning
We use the Unsloth framework [11] and apply a consistent set of hyperparameters for the three fine-tuning settings: without tuning, synthetic tuning, and mixed tuning, as shown in Table 7. The fine-tuning process is performed using a single NVIDIA A100-SXM4-40GB GPU.
Category Detail
Model model_name = “unsloth/Qwen2-VL-7B-Instruct”, load_in_4bit = True
LoRA r = 16, lora_alpha = 16, lora_dropout = 0, random_state = 3407,
Fine-tuning per_device_train_batch_size = 2, gradient_accumulation_steps = 4, warmup_steps = 5, max_steps = 30, learning_rate = 2e-4,
optim = “adamw_8bit”, weight_decay = 0.01, lr_scheduler_type = “linear”,
seed = 3407, max_seq_length = 2048,
Table 7: Fine-tuning Configuration and Hyperparameters.
C.6 Performance of Selected Models on Existing Benchmarks
We obtained the accuracy of all selected models on the four existing datasets from the OpenVLM Leaderboard [^5] https://huggingface.co/spaces/opencompass/open_vlm_leaderboard. To ensure consistency with the evaluation metric used in GuessBench, which is the average accuracy per question, we selected the Overall Accuracy metric for the AI2D, MMBench V1.1 Test (EN), and MathVista benchmarks. For HallusionBench, we instead used aAcc, which represents the overall accuracy across all atomic questions.
Dataset GPT-4o Gemini-2.0-Flash InternVL2.5-78B-MPO Qwen2.5VL-72B MiniCPM-V2.6
AI2D 84.9 83.1 89.2 88.5 82.1
MMBench V1.1 Test(EN) 84.8 70.4 87.9 88.3 79.0
MathVista 60.0 70.4 76.6 74.2 60.8
HallusionBench 71.4 72.0 73.3 71.9 65.0
Table 8: The detailed accuracy of all models on the four existing datasets.
Appendix D Sycophantic Behavior in VLMs under Contradictory User Feedback
To investigate whether VLMs exhibit sycophantic behavior when confronted with contradictory user requests, we conduct a study focusing on two state-of-the-art models: GPT-4o and Gemini-2.0-Flash. Within the static evaluation setting, we select 100 instances from the Minecraft build sets that both models originally answered correctly. For each instance, after the VLM provides a correct response, we intentionally assert that its answer is incorrect, allowing us to observe whether the model revises its correct prediction and instead conforms to the user’s false feedback.
We design three conditions to induce such behavior in VLMs. In the Base setting, we merely inform the VLM that its answer is wrong without providing any additional information. The Random setting builds upon the Base condition by additionally supplying a randomly chosen incorrect answer, sampled from the set of incorrect options in GuessBench, and asserting it as correct. The Similar setting also builds upon the Base condition, but the misleading answer is semantically similar to the correct one, generated using GPT-4o, and presented as the correct answer.
The results reveal that GPT-4o maintains an accuracy of 6%, 15%, and 26% under the Base, Random, and Similar settings respectively. In contrast, Gemini-2.0-Flash achieves only 1%, 2%, and 4% accuracy across the same conditions. These findings suggest that Gemini-2.0-Flash is more susceptible to sycophantic behavior than GPT-4o.
Interestingly, VLMs are most likely to revise their initially correct answers in the Base setting, where no alternative answer is proposed, and are least likely to do so in the Similar setting, where a plausible but incorrect alternative is given. Further analysis of the model responses under contradictory user feedback reveals that in both the Random and Similar settings, the VLMs do not engage in any substantive evaluation of the user-supplied answer’s plausibility. Instead, they often respond with generic phrases such as “reconsidering the build and the hint,” immediately followed by a new, often incorrect, guess.
Paradoxically, as the misleading input becomes more semantically aligned with the correct answer from Base to Random to Similar, the models’ accuracy increases. This trend runs counter to our hypothesis. If the models were capable of independent reasoning, misleading cues that closely resemble the correct answer would be more deceptive and thus reduce accuracy. Instead, our results suggest that VLMs fundamentally lack the ability to assess correctness or engage in independent thought. Rather, they merely generate text that appears syntactically plausible in response to user prompts.
Appendix E Flipping the Paradigm: Easier to Generate than Discriminate
Existing experiments show that generative tasks are generally more challenging for LMs than discriminative ones [29]. However, our findings reveal a surprising result: creatively generating a Minecraft build is significantly easier than identifying one, further underscoring the novelty and distinctiveness of GuessBench.
In our generative experiment setup, we select the Gemini-2.0-Flash model and randomly sample 100 build sets from GuessBench. For each sampled item, we manually input prompts into the Gemini playground [^6] https://aistudio.google.com/prompts/new_chat, asking the model to generate an image of a Minecraft build based on a given answer, thereby reversing the question-answering direction in GuessBench. For the discriminative counterpart, we adopt the static task from GuessBench as our evaluation setting.
We evaluate performance using accuracy. A generation is considered correct only if it satisfies two criteria: it contains the specified entities, and it appears plausibly constructed within the Minecraft environment. Under this rigorous standard, Gemini-2.0-Flash achieves an accuracy of 80% on the generative task, which is substantially higher than its 40% accuracy on the corresponding discriminative task within GuessBench.
This notable performance gap highlights a critical limitation in VLMs’ ability to perform creative understanding in open-ended scenarios, a challenge that remains largely overlooked in prior work. GuessBench not only exposes this limitation but also points to promising directions for future improvements in VLM capabilities.
Appendix F From Feedback to Foresight: GPT-4o Learns in Context Like a Player
To more accurately simulate the gameplay of “Guess the Build” and evaluate the performance of VLMs in a realistic gaming environment, we divide a total of 500 questions into 100 sets, each containing 5 questions. This setup mirrors the actual game mechanics, where five questions are asked consecutively, and after each question, the VLM receives feedback indicating whether its guess is correct or incorrect. In our experiments, we adopt GPT-4o as the VLM under investigation.
As shown in Figure 8, answering five questions in succession leads to improved accuracy, with a more substantial performance gain observed in dynamic tasks. Specifically, accuracy increases by 4.8% in static tasks and by an average of 10.47% across three attempts in dynamic tasks. These results indicate that GPT-4o is able to extract more information from dynamic tasks and utilize prior interactions through in-context learning to enhance its responses to subsequent questions. Moreover, we observe that in dynamic tasks, the improvement in accuracy diminishes with each additional attempt. This suggests that the benefits of GPT-4o’s in-context learning are more pronounced when dealing with more incomplete Minecraft builds and less informative hints.
This experiment demonstrates that GPT-4o possesses strong in-context learning capabilities, enabling it to answer new questions more effectively by drawing on previous ones, similar to human players. It also highlights a promising direction for enhancing VLMs’ ability to understand creativity in real-world contexts.
Figure 8: Evaluation results of Single-Question and Multi-Question settings. All denotes the overall accuracy for each task. Answering multiple questions sequentially within the same context improves response accuracy.
Appendix G Reasoning Reliability Analysis
To evaluate whether VLMs can accurately explain their reasoning path when correctly identifying a Minecraft build, we conduct a manual analysis under the static task setting. Specifically, we randomly sample 50 responses from all correctly answered questions by GPT-4o and examine the reasoning path provided in each response. We find that 14% of the responses contain issues in their reasoning paths. Among these, 42.8% contain flawed reasoning (for example, inferring that a build resembles “a pit or a mine with ores inside” based on appearance, but ultimately identifying it as a “crate”), 28.6% involve errors in visual perception (such as stating that a table has four legs when the image clearly shows only three), and the remaining 28.6% result from inaccurate textual perception (for instance, misinterpreting the hint “_ l a _ _ _ _” as “l a”). These findings suggest that GPT-4o occasionally generates incorrect reasoning even when the final answer is accurate, underscoring its limitations in creativity understanding.
Appendix H Examples
H.1 Illustrative Examples for Static and Dynamic Tasks
For both the static and dynamic task settings, we selected the example “Graveyard” to illustrate the entire process within each task, including how GPT-4o receives the question, produces a response, and is subsequently evaluated for correctness. This example is intended to help readers gain a clearer understanding of the specific nature and structure of the two tasks.
Figure 9: An example illustrates the first attempt in the dynamic task setting: GPT-4o responds based on an incomplete Minecraft build and a partial hint, but produces an incorrect answer.
Figure 10: An example illustrates the second attempt in the dynamic task setting: GPT-4o, with access to the details of the first attempt, responds again based on a more complete Minecraft build and a more informative hint, but still produces an incorrect answer.
Figure 11: An example illustrates the third attempt in the dynamic task setting: GPT-4o, with access to the details of the previous two attempts, responds again based on the most complete Minecraft build and the most comprehensive hint, and produces a correct answer.
Figure 12: An example of the static task setting: GPT-4o, based on a complete Minecraft build and a full hint, makes a prediction and provides the correct answer.
H.2 Case Study: Image Retrieval Reduces Answer Accuracy
We present a case where the use of an Image Retrieval method unexpectedly reduces the accuracy of the model’s prediction. Initially, GPT-4o correctly predicts the answer as Modern Art. However, since Modern Art encompasses a wide range of architectural styles and painting techniques, the image retrieval method returns an image of Starry Night by Vincent van Gogh. Although this painting falls under the category of Modern Art, it differs significantly from the original Minecraft build in visual appearance. As a result, GPT-4o alters its originally correct prediction. Subsequent image retrievals and comparisons fail to guide the model back to the correct answer.
Figure 13: One example from the static task. Initially, GPT-4o provides the correct answer based on the Minecraft build and the accompanying hint.
Figure 14: Following the results of the first image retrieval, GPT-4o revises its initially accurate prediction after retrieving and comparing the images, resulting in an incorrect response. At this stage, GPT-4o remains in the refinement process and has not yet concluded, as it has not undergone evaluation.
Figure 15: Subsequently, GPT-4o performs a second round of image retrieval. After retrieving and comparing the images, it again modifies its original prediction; however, the response remains incorrect. At this point, GPT-4o is still in the refinement process and has not yet reached completion, as it has not undergone final evaluation.
Figure 16: Finally, GPT-4o conducts a third round of image retrieval. After retrieving and comparing the images, it concludes that its original prediction is correct. It therefore submits “Finger of God”, proposed in the previous round, as its final answer and proceeds to evaluation. However, this answer is incorrect.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Please refer to the abstract and §1 for more details.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Please refer to the Appendix A for more details.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Please refer to the §3 and Appendix C for more details.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of API models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Our code and data are available at https://anonymous.4open.science/r/GuessBench-E863/.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Please refer to the §3 and Appendix C for more details.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [N/A]
  34. Justification: We do not propose a new method or algorithm, nor do we claim it outperforms baselines; therefore, we do not conduct statistical significance tests.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Please refer to the §3 and Appendix C for more details.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Please refer to the §2.4 and Appendix B for more details.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Please refer to the §2.4 and Appendix B for more details.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [N/A]
  59. Justification: We use open data and models allowed for academic research.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Our code and data are available at https://anonymous.4open.science/r/GuessBench-E863/, and the limitations are discussed in Appendix A.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLM is used solely for grammar checking and language polishing during writing.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.